OpenStudy (anonymous):

some one hewp plzzzzz Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

thank you you were alot of help to me :)

OpenStudy (opcode):

Okay first we got to understand your problem: \(\text{Create a unique example of dividing a polynomial by a monomial and provide}\) \(\text{the simplified form. }\) So let's set some questions up: \(1) x^2 + 8x + 12 / x+ 2\) \(2)x - 2 +1/3x\) Okay for number one follow the steps below: 1) Factor out the numerator, so now you have \((x+2)(x+6).\) 2) Cancel out \((x+2)\) in the numerator and in the detonator. 3) Now you are left with the simplified answer, which is \(x+ 6\). Okay for number two follow the steps below: 1) Multiply \(x - 2 +1/3x\) by \(3x^2\) 2) You get \(3x^3 - 2x^2 + 3x^2/ 3x\) 3) That simplifies to x \(\text{Explain, in complete sentences, the two ways used to simplify this expression}\) \(\text{and how you would check your quotient for accuracy.}\) What I did: I simplified the expression. I explained how I simplified the expression. What you do: Check your quotient for accuracy. @mikala1 I believe you know how to do that? (If you don't I'll do it for you.)

Parth (parthkohli):

I like the first one, but the second one is a little wrong. :-)

OpenStudy (opcode):

Tell me how to fix it. There's a medal in it for you ^_^.

OpenStudy (opcode):

Oh. Such a stupid mistake I made. Now I feel silly.

Parth (parthkohli):

Another way:\[\begin{aligned}= & \dfrac{x -1}{3x} \\ \\ \\ = & \dfrac{x}{3x} - \dfrac{1}{3x} \\ \\ \\ = & \dfrac{1}{3} - \dfrac{1}{3x} \end{aligned}\]I did a mistake in the last post too :\

OpenStudy (opcode):

It's all good. Thank you Parth. I just don't know why I forgot such a vital step :| It's sleep deprivation. I've even answered History questions incorrectly :S

Parth (parthkohli):

Nah, happens, like it happened to me above.